# Diagnostic-Guided Dynamic Profile Optimization for LLM-based User Simulators in Sequential Recommendation

### v3

@@ Abstract
Recent advances in large language models (LLMs) have enabled realistic user simulators for developing and evaluating recommender systems (RSs). However, existing LLM-based simulators for RSs face two major limitations: (1) static and single-step prompt-based inference that leads to inaccurate and incomplete user profile construction; (2) unrealistic and single-round recommendation-feedback interaction pattern that fails to capture real-world scenarios. To address these limitations, we propose DGDPO (Diagnostic-Guided Dynamic Profile Optimization), a novel framework that constructs user profile through a dynamic and iterative optimization process to enhance the simulation fidelity. Specifically, DGDPO incorporates two core modules within each optimization loop: firstly, a specialized LLM-based diagnostic module, calibrated through our novel training strategy, accurately identifies specific defects in the user profile. Subsequently, a generalized LLM-based treatment module analyzes the diagnosed defect and generates targeted suggestions to refine the profile. Furthermore, unlike existing LLM-based user simulators that are limited to single-round interactions, we are the first to integrate DGDPO with sequential recommenders, enabling a bidirectional evolution where user profiles and recommendation strategies adapt to each other over multi-round interactions. Extensive experiments conducted on three real-world datasets demonstrate the effectiveness of our proposed framework. 00footnotetext: denotes the corresponding authors.
@@ Introduction
User simulators (Balog and Zhai 2023), which aim to mimic user behaviors and decision-making patterns, are crucial for advancing recommender systems (RSs) (Luo et al. 2022). They offer a controllable and cost-effective paradigm for developing and evaluating RSs, reducing the reliance on real-user experiments that are often expensive and ethically sensitive. While early user simulators ranged from rule-based (Rohde et al. 2018) to deep learning-based approaches (Shi et al. 2019b; Zhao et al. 2021), these traditional methods fail to explicitly represent users’ complex semantic preferences and generate diverse user interaction behaviors. To address these issues, Large Language Models (LLMs) have recently emerged as a new paradigm (Sun et al. 2024b). With their advanced cognitive and generative capabilities, LLMs are typically integrated into an agent framework (Park et al. 2023) that incorporates components such as Profile, Memory, and Action to emulate user interactions with RSs. For example, Agent4Rec (Zhang et al. 2024a) and RecAgent (Wang et al. 2025) prompt LLMs to generate user profile, enabling them to simulate diverse behaviors that mirror the complexity of real users.
However, existing LLM-based user simulators face two major limitations: (1) the inaccuracy and incompleteness of profile construction due to the static and single-step prompt-based inference. By generating the user profile only at initialization, user simulators produce a static representation that cannot adapt to the evolving user interests across multi-round interactions. Thus, any inaccuracy or incompleteness in this initial profile persists uncorrected, causing the simulated behavior to progressively diverge from real user actions throughout the simulation; (2) the unrealistic and single-round recommendation-feedback interaction pattern. Most existing simulators are typically integrated with static recommenders, e.g., MF (Koren, Bell, and Volinsky 2009), failing to simulate authentic, multi-round interactions where user profile and RS strategies should adapt and evolve based on the latest feedback.
@@
Figure 1: Accuracy comparison for profile defect identification. Our specialized LLM-based diagnostic module significantly outperforms the prompt-based general-purpose LLM. Details are included in Appendix D.
Therefore, we propose a novel Diagnostic-Guided Dynamic Profile Optimization (DGDPO) framework for a more realistic simulation. However, developing this framework faces two primary challenges. (CH1): How to build a reliable mechanism for dynamic profile construction? While existing methods (Shu et al. 2024) like self-reflection (Sun et al. 2024a) attempt iterative refinement, they often rely on general-purpose LLMs, thus may not be reliable due to potential hallucinations. As shown in Fig. 1, directly prompting general-purpose LLMs only achieves an average accuracy of 62.78% on profile defect identification. This highlights the necessity of a more targeted strategy that can accurately diagnose the specific type of defect, such as inaccurate and/or incomplete, and then apply a tailored refinement. To achieve this, we first design a specialized LLM-based diagnostic module through domain-adaptive pre-training and defect-specific fine-tuning, enabling reliable identification of profile defects. As shown in Fig. 1, the module achieves an average accuracy of 92.20%, significantly outperforming the general-purpose LLM. We then design a generalized LLM-based treatment module that uses this precise diagnosis to analyze the defect and generate targeted refinements for the profile. Instead of single-step prompting, our framework iteratively optimizes the profile by traversing the user’s interaction history step-by-step, to dynamically enhance its accuracy and comprehensiveness. (CH2): How to achieve realistic multi-round interactions where user profiles and recommender strategies mutually influence and evolve over time? Inspired by the multi-round interaction nature of real-world recommendation scenarios, we propose to incorporate the user simulator with sequential recommenders (SRs) (Sun et al. 2025). This establishes a genuine bidirectional evolution: the user profile is updated based on interactions, while the SRs adaptively adjust their recommendation strategies in response to the evolving user behavior.
@@
Traditional User Simulators for RSs. Early user simulators for RSs progressed from rule-based approaches (Rohde et al. 2018; Ie et al. 2019; Shi et al. 2019a) to advanced data-driven techniques (Huang et al. 2020), including methods based on reinforcement learning (RL) (Shi et al. 2019b; Zou et al. 2020), generative adversarial networks (GANs) (Chen et al. 2019; Shi et al. 2019b; Zhao et al. 2021), and transformers (Zhao et al. 2023; Afzali et al. 2023). However, they lack external context and knowledge, as well as the sophisticated cognitive reasoning mechanism that real users leverage during decision-making processes.
LLM-based User Simulators for RSs. The advent of LLMs has introduced novel paradigms for user simulation in RSs (Cai et al. 2025; Ma et al. 2025). LLM-based simulators have been widely adopted across various applications, including evaluating multi-round conversational systems (Friedman et al. 2023; Zhu, Huang, and Sang 2024) and generating high-quality interaction data for RL-based algorithms (Corecco et al. 2024; Ebrat, Paradalis, and Rueda 2024; Zhang et al. 2025). Current works on user profile construction can be categorized into two types. Some works (Zhang et al. 2024a; Wang et al. 2025) create user profile once at initialization, resulting in static representations that cannot capture evolving user preferences. Others employ self-reflection (Zhang et al. 2024b; Shu et al. 2024) for iterative refinement. However, they suffer from two limitations: (1) they adopt general-purpose LLMs which may generate unreliable profile due to potential hallucinations; (2) they rely on unrealistic single-round interactions, lacking dynamic adaptation to user feedback and failing to simulate the bidirectional evolution between user profile and recommendation strategies in realistic interactions.
Sequential Recommenders. SRs (Du et al. 2024) have evolved from early approaches combining Markov Chains with matrix factorization (Rendle, Freudenthaler, and Schmidt-Thieme 2010), to deep learning models. These include CNNs for local pattern extraction (Tang and Wang 2018), RNNs to capture temporal dependencies (Hidasi 2015; Huang et al. 2025), attention mechanisms to focus on relevant historical items (Li et al. 2017; Kang and McAuley 2018; Liu et al. 2018; Sun et al. 2019), and GNNs to model high-order relationships within sequential data (Wu et al. 2022; Yu et al. 2024; Li et al. 2024; Liu, Xia, and Huang 2024). Despite their success, they have primarily been evaluated in a static mode, which may fail to fully reflect the dynamic complexity of real-world user interactions.
@@
Given a user’s historical interaction sequence $\mathcal{D}_{u}=[i_{1},i_{2},...,i_{T}]$ ordered by timestamp, we split it into two portions: $\mathcal{D}_{u}^{(ini)}=[i_{1},i_{2},...,i_{n}]$, and $\mathcal{D}_{u}^{(opt)}=[i_{n+1},...,i_{T}]$. For effective and efficient user profile construction, we first create an initial user profile $\mathcal{P}_{u}^{(ini)}$ for each simulated user $u$. It is derived from $\mathcal{D}_{u}^{(ini)}$ via a single-step prompt-based inference method (Zhang et al. 2024a), serving as the foundation user profile. Then, we leverage $\mathcal{D}_{u}^{(opt)}$ to dynamically optimize $\mathcal{P}_{u}^{(ini)}$ via our proposed DGDPO, ultimately obtaining an optimized user profile $\mathcal{P}_{u}^{(opt)}$.
@@
To construct the pre-training corpus, for each user $u$, we leverage his initial sequence data $\mathcal{D}_{u}^{(ini)}$. For each item $i_{t}\in\mathcal{D}_{u}^{(ini)}$, we instruct the user simulator with profile $\mathcal{P}_{u}^{(ini)}$ to make a binary interaction decision for $i_{t}$ and generate a textual rationale. To ensure the data quality, we only consider the case where the user simulator’s decision matches with the real user $u$’s behavior (i.e., choosing to interact with $i_{t}$) and user $u$ gave item $i_{t}$ a high rating ($\geq$ 3 out of 5). Then, the user simulator’s output, together with $\mathcal{P}_{u}^{(ini)}$ and relevant information (e.g., title, genres, and rating) of $i_{t}$, is used to construct a comprehensive corpus of structured texts, where the details are shown in Appendix A.
@@
The pipeline starts with the high-quality, non-defective profile-item pairs that constitute our pre-training corpus. Building upon these non-defective pairs, we subsequently perform a series of targeted modifications to generate synthetic user profiles with different defect types: (1) Inaccurate: to create a contradiction, we first identify preference descriptions in the profile relevant to the target item (e.g., its genre). We then generate a new description with the opposite sentiment (e.g., transforming “enjoys comedy films” to “dislikes comedy films”) and use it to replace the original statement. (2) Incomplete: to simulate a lack of information, we randomly remove a proportion of descriptions relevant to the target item from the profile, making the profile insufficient to explain the user’s interaction. (3) Inaccurate & Incomplete: we first apply the “incomplete” modification, followed by the “inaccurate” modification. Using these synthetic profiles, we then format the data for instruction fine-tuning, where each sample has four parts: System, Instruction, Input, and Output, as shown in Appendix A.
@@
  // Single-step prompt-based inference
@@
         // The Specialized LLM-based Diagnostic Module
@@ 5       $\{l\}\leftarrow f_{\text{diag}}(\mathcal{P}_{u}^{(ini)},\mathcal{C})$;
         // The Generalized LLM-based Treatment Module
       $\{r\}\leftarrow\text{Infer}(\mathcal{P}_{u}^{(ini)},\mathcal{C},\{l\})$;
          // Infer defect reasons
         // Generate optimization suggestions
@@ 6       $\mathcal{O}\leftarrow\text{GenOpt}(\mathcal{P}_{u}^{(ini)},\mathcal{C},\{l\},\{r\})$;
       $\tilde{\mathcal{P}}_{u}^{(ini)}\leftarrow\text{Refine}(\mathcal{P}_{u}^{(ini)},\mathcal{O})$;
          // Refine the profile
7       $\mathcal{C}$ $\leftarrow$ [ ];
@@ 8       $\mathcal{P}_{u}^{(ini)}$ $\leftarrow$ $\tilde{\mathcal{P}}_{u}^{(ini)}$;
This module is responsible for the complex reasoning and modification tasks. While the specialized diagnostic module excels at identifying what defects exist within a profile, understanding why these defects occur and determining how to address them effectively requires sophisticated common-sense reasoning and high-fidelity generative capabilities. Consequently, we employ a powerful general-purpose LLM, leveraging its superior analytical reasoning and text generation capabilities to carry out the following treatment process [^1] Detailed prompts are attached in Appendix B..
@@
Existing LLM-based user simulators are primarily integrated with static recommenders in page-by-page scenarios (Zhang et al. 2024a), which essentially represent a one-time recommendation result split into multiple pages. This fails to capture realistic scenarios where users continuously interact with the RSs, leading to the dynamic evolution of both user profiles and recommendation strategies through user feedback and interactions. To enable realistic multi-round interactions, we propose to incorporate our optimized user simulator DGDPO with sequential recommenders (SRs). Unlike traditional static models, SRs excel at capturing temporal dynamics, which makes them ideal for modeling the bidirectional evolution of user preferences and recommendation strategies over time.
@@
We conduct extensive experiments to answer five research questions. (RQ1) Can DGDPO improve the fidelity of user simulators? (RQ2) How do different profile update strategies and SRs impact the simulation fidelity in multi-round interaction scenarios? (RQ3) How do various components of DGDPO affect its performance? (RQ4) How do essential hyper-parameters affect DGDPO? (RQ5) How does DGDPO construct user profile with its optimization process?
@@
We use three real-world datasets from various domains. Specifically, ML-1M (Harper and Konstan 2015) is widely used in movie recommendations. Amazon Books (Books) and Amazon Movies & TV (Movies & TV) are product review datasets collected from Amazon for Books and Movies & TV categories (Ni, Li, and McAuley 2019), respectively. We filter out users and items with less than 5 interactions in ML-1M, and those with less than 20 interactions in Amazon datasets. To support multi-round interactions, we first sort each user’s interaction chronologically. We set the maximum sequence length to 200 for ML-1M and 50 for Amazon datasets. For each user $u$, we select the most recent 10 items as the test set, the 11th most recent item as the validation set, and the remaining as the training set (i.e., $\mathcal{D}_{u}$). Particularly, we further split $\mathcal{D}_{u}$ into $\mathcal{D}_{u}^{(ini)}$ and $\mathcal{D}_{u}^{(opt)}$ with the ratio of $\alpha:(1-\alpha)$. The statistics of these datasets after preprocessing are summarized in Table 1.
@@
We compare DGDPO with five baselines for user profile construction in two categories. In traditional methods, KuaiSim (Zhao et al. 2023) is a transformer-based model for user profile construction. In LLM-based methods, Agent4Rec (Zhang et al. 2024a) utilizes a single-step prompt-based inference. LLM-US (Zhang et al. 2025) generates explanations of user ratings as profile. RAH (Shu et al. 2024) and AgentCF (Zhang et al. 2024b) employ self-reflection and collaborative reflection, respectively. To enable multi-round interactions, we integrate the user simulator with four representative SRs. The conventional method FPMC (Rendle, Freudenthaler, and Schmidt-Thieme 2010) combines matrix factorization with first-order Markov chains. The deep learning-based methods include NARM (Li et al. 2017), an RNN-based model with an attention mechanism; SASRec (Kang and McAuley 2018), using unidirectional self-attention; and LRURec (Yue et al. 2024), leveraging a linear recurrent unit-based model.
@@
Following state-of-the-arts (Shu et al. 2024; Zhang et al. 2024a), we randomly sample 1000 users from each dataset to serve as simulated users. Our diagnostic module is based on lightweight Qwen3-0.6B (Yang et al. 2025), while the treatment module employs gpt-4o-mini via its API with a temperature of 0. All LLM-based baselines also utilize GPT-4o-mini as their backbone. For DGDPO, we set $N_{1}=4$, $\alpha=0.6$, and $\mathcal{X}=20$ for all datasets. We build upon Agent4Rec (Zhang et al. 2024a) as the foundation user simulator and conduct ten-round interactions for each user. For KuaiSim and SRs methods, we tune them by Optuna (optuna.org) with 50 trials and train for up to 100 epochs with an early stopping mechanism (Sun et al. 2022). All experiments are conducted on an NVIDIA V100 GPU. More detailed implementations are in Appendix C.
@@
Following (Luo et al. 2022; Zhang et al. 2024a), we adopt Precision (Prec), Recall, Accuracy (Acc), and F1 Score (F1) as the evaluation metrics to measure the consistency between the simulated and real users.
@@
Fig. 3 depicts the performance of DGDPO across ten interaction rounds on Books dataset under different profile update strategies [^2] Hereafter, we only show the results on Books due to space limitation, and similar trends can be noted on the other two datasets.. The results unveils that dynamic profile update in multi-round interactions can further improve the simualtion accuracy. (1) For ‘w/o Update’, the performance generally exhibits minimal fluctuations across multi-round interactions and is the worst among all strategies. With profile unchanged, ‘w/o Update’ fails to capture the evolution of user preference during multi-round interactions, thereby limiting the simulator’s long-term fidelity. (2)Update w/o GT’ exhibits performance divergence when interacting with different SRs. Specifically, with FPMC, the performance shows a declining trend. In contrast, with advanced deep learning-based SRs, we observe steady performance improvements. This occurs because advanced SRs rank ground truth items higher, and LLM-based user simulators, sensitive to item positions (Hou et al. 2024), are more likely to interact with them when they are prominently placed, resulting in more accurate simulation. (3)Update w/ GT’ performs the best among all methods. This is attributed to an idealized update strategy that always uses the ground truth item to update the user profile, regardless of the simulator’s actions. However, in practical applications, ground truth items are often unavailable, which limits the applicability of this strategy. Therefore, it underscores the importance of developing more advanced user simulators that closely replicate real user behaviors, enabling more effective and reliable updates to the user profile during long-term interactions.
@@
Fig. 5 provides a concrete illustration of DGDPO’s optimization process on Books dataset. It is triggered by observing a discrepancy case where the simulator built upon $\mathcal{P}_{u}^{(ini)}$ did not choose to interact with the book “Munich, 1938: Appeasement and World War II”, while the real user read it and gave it a high rating of 5. At Step 2, our diagnostic module then correctly identifies this gap and diagnoses the profile as “Incomplete” (in red). Guided by this diagnosis, at Step 3, the treatment module infers the underlying reason (in yellow), and generates a targeted suggestion to add the missing preference for “Politics & Social Sciences” genres (in green) at Step 4. At Step 5, this suggestion is applied to obtain the updated profile $\tilde{\mathcal{P}}_{u}^{(ini)}$ (in blue), making it more comprehensive and capable of explaining the user’s interest in the book. Additional case studies regarding “Inaccurate” and “Inaccurate & Incomplete” are in Appendix E.

### v4

@@ Abstract
Recent advances in large language models (LLMs) have enabled realistic user simulators for developing and evaluating recommender systems (RSs). However, existing LLM-based simulators for RSs face two major limitations: (1) static and single-step prompt-based inference that leads to inaccurate and incomplete user profile construction; (2) unrealistic and single-round recommendation-feedback interaction pattern that fails to capture real-world scenarios. To address these limitations, we propose DGDPO (Diagnostic-Guided Dynamic Profile Optimization), a novel framework that constructs user profile through a dynamic and iterative optimization process to enhance the simulation fidelity. Specifically, DGDPO incorporates two core modules within each optimization loop: firstly, a specialized LLM-based diagnostic module, calibrated through our novel training strategy, accurately identifies specific defects in the user profile. Subsequently, a generalized LLM-based treatment module analyzes the diagnosed defect and generates targeted suggestions to refine the profile. Furthermore, unlike existing LLM-based user simulators that are limited to single-round interactions, we are the first to integrate DGDPO with sequential recommenders, enabling a bidirectional evolution where user profiles and recommendation strategies adapt to each other over multi-round interactions. Extensive experiments conducted on three real-world datasets demonstrate the effectiveness of our proposed framework.
@@ Introduction
User simulators (Balog:2024:FnTIR), which aim to mimic user behaviors and decision-making patterns, are crucial for advancing recommender systems (RSs) (luo2022mindsim). They offer a controllable and cost-effective paradigm for developing and evaluating RSs, reducing the reliance on real-user experiments that are often expensive and ethically sensitive. While early user simulators ranged from rule-based (rohde2018recogym) to deep learning-based approaches (shi2019virtual; zhao2021usersim), these traditional methods fail to explicitly represent users’ complex semantic preferences and generate diverse user interaction behaviors. To address these issues, Large Language Models (LLMs) have recently emerged as a new paradigm. With their advanced cognitive and generative capabilities, LLMs are typically integrated into an agent framework (park2023generative) that incorporates components such as Profile, Memory, and Action to emulate user interactions with RSs. For example, Agent4Rec (zhang2024generative) and RecAgent (wang2025user) prompt LLMs to generate user profile, enabling them to simulate diverse behaviors that mirror the complexity of real users.
However, existing LLM-based user simulators face two major limitations: (1) the inaccuracy and incompleteness of profile construction due to the static and single-step prompt-based inference. By generating the user profile only at initialization, user simulators produce a static representation that cannot adapt to the evolving user interests across multi-round interactions. Thus, any inaccuracy or incompleteness in this initial profile persists uncorrected, causing the simulated behavior to progressively diverge from real user actions throughout the simulation; (2) the unrealistic and single-round recommendation-feedback interaction pattern. Most existing simulators are typically integrated with static recommenders, e.g., MF (koren2009matrix), failing to simulate authentic, multi-round interactions where user profile and RS strategies should adapt and evolve based on the latest feedback.
@@
Figure 1: Accuracy comparison for profile defect identification. Our specialized LLM-based diagnostic module significantly outperforms the prompt-based general-purpose LLM.
Therefore, we propose a novel Diagnostic-Guided Dynamic Profile Optimization (DGDPO) framework for a more realistic simulation. However, developing this framework faces two primary challenges. (CH1): How to build a reliable mechanism for dynamic profile construction? While existing methods (shu2024rah) like self-reflection (sun2024large) attempt iterative refinement, they often rely on general-purpose LLMs, thus may not be reliable due to potential hallucinations. As shown in Fig. 1, directly prompting general-purpose LLMs only achieves an average accuracy of 62.78% on profile defect identification. This highlights the necessity of a more targeted strategy that can accurately diagnose the specific type of defect, such as inaccurate and/or incomplete, and then apply a tailored refinement. To achieve this, we first design a specialized LLM-based diagnostic module through domain-adaptive pre-training and defect-specific fine-tuning, enabling reliable identification of profile defects. As shown in Fig. 1, the module achieves an average accuracy of 92.20%, significantly outperforming the general-purpose LLM. We then design a generalized LLM-based treatment module that uses this precise diagnosis to analyze the defect and generate targeted refinements for the profile. Instead of single-step prompting, our framework iteratively optimizes the profile by traversing the user’s interaction history step-by-step, to dynamically enhance its accuracy and comprehensiveness. (CH2): How to achieve realistic multi-round interactions where user profiles and recommender strategies mutually influence and evolve over time? Inspired by the multi-round interaction nature of real-world recommendation scenarios, we propose to incorporate the user simulator with sequential recommenders (SRs) (sun2025llm4rsr). This establishes a genuine bidirectional evolution: the user profile is updated based on interactions, while the SRs adaptively adjust their recommendation strategies in response to the evolving user behavior.
@@
Traditional User Simulators for RSs. Early user simulators for RSs progressed from rule-based approaches (rohde2018recogym; ie2019recsim; shi2019pyrecgym) to advanced data-driven techniques (huang2020keeping), including methods based on reinforcement learning (RL) (shi2019virtual; zou2020pseudo), generative adversarial networks (GANs) (chen2019generative; shi2019virtual; zhao2021usersim), and transformers (zhao2023kuaisim; afzali2023usersimcrs). However, they lack external context and knowledge, as well as the sophisticated cognitive reasoning mechanism that real users leverage during decision-making processes.
LLM-based User Simulators for RSs. The advent of LLMs has introduced novel paradigms for user simulation in RSs (cai2025agentic; ma2025pub; wei2025mirroring). LLM-based simulators have been widely adopted across various applications, including evaluating conversational systems (friedman2023leveraging; zhu2024reliable; zhu2025llm) and generating high-quality interaction data for RL-based algorithms (corecco2024llm; ebrat2024lusifer; zhang2025llm). Current works on user profile construction can be categorized into two types. Some works (zhang2024generative; wang2025user) create user profile once at initialization, resulting in static representations that cannot capture evolving user preferences. Others employ self-reflection (zhang2024agentcf; shu2024rah) for iterative refinement. However, they suffer from two limitations: (1) they adopt general-purpose LLMs which may generate unreliable profile due to potential hallucinations; (2) they rely on unrealistic single-round interactions, lacking dynamic adaptation to user feedback and failing to simulate the bidirectional evolution between user profile and recommendation strategies in realistic interactions.
Sequential Recommenders. SRs have evolved from early approaches combining Markov Chains with matrix factorization (rendle2010factorizing), to deep learning models. These include CNNs for local pattern extraction (tang2018personalized), RNNs to capture temporal dependencies (hidasi2015session; zhu2017next; zhu2019query), attention mechanisms to focus on relevant historical items (li2017neural; kang2018self; liu2018stamp; sun2019bert4rec), and GNNs to model high-order relationships within sequential data (wu2022gcrec; yu2024learning; li2024global; liu2024selfgnn). Despite their success, they have primarily been evaluated in a static mode, which may fail to fully reflect the dynamic complexity of real-world user interactions.
@@
Given a user’s historical interaction sequence $\mathcal{D}_{u}=[i_{1},i_{2},...,i_{T}]$ ordered by timestamp, we split it into two portions: $\mathcal{D}_{u}^{(ini)}=[i_{1},i_{2},...,i_{n}]$, and $\mathcal{D}_{u}^{(opt)}=[i_{n+1},...,i_{T}]$. For effective and efficient user profile construction, we first create an initial user profile $\mathcal{P}_{u}^{(ini)}$ for each simulated user $u$. It is derived from $\mathcal{D}_{u}^{(ini)}$ via a single-step prompt-based inference method (zhang2024generative), serving as the foundation user profile. Then, we leverage $\mathcal{D}_{u}^{(opt)}$ to dynamically optimize $\mathcal{P}_{u}^{(ini)}$ via our proposed DGDPO, ultimately obtaining an optimized user profile $\mathcal{P}_{u}^{(opt)}$.
@@
To construct the pre-training corpus, for each user $u$, we leverage his initial sequence data $\mathcal{D}_{u}^{(ini)}$. For each item $i_{t}\in\mathcal{D}_{u}^{(ini)}$, we instruct the user simulator with profile $\mathcal{P}_{u}^{(ini)}$ to make a binary interaction decision for $i_{t}$ and generate a textual rationale. To ensure the data quality, we only consider the case where the user simulator’s decision matches with the real user $u$’s behavior (i.e., choosing to interact with $i_{t}$) and user $u$ gave item $i_{t}$ a high rating ($\geq$ 3 out of 5). Then, the user simulator’s output, together with $\mathcal{P}_{u}^{(ini)}$ and relevant information (e.g., title, genres, and rating) of $i_{t}$, is used to construct a comprehensive corpus of structured texts.
@@
The pipeline starts with the high-quality, non-defective profile-item pairs that constitute our pre-training corpus. Building upon these non-defective pairs, we subsequently perform a series of targeted modifications to generate synthetic user profiles with different defect types: (1) Inaccurate: to create a contradiction, we first identify preference descriptions in the profile relevant to the target item (e.g., its genre). We then generate a new description with the opposite sentiment (e.g., transforming “enjoys comedy films” to “dislikes comedy films”) and use it to replace the original statement. (2) Incomplete: to simulate a lack of information, we randomly remove a proportion of descriptions relevant to the target item from the profile, making the profile insufficient to explain the user’s interaction. (3) Inaccurate & Incomplete: we first apply the “incomplete” modification, followed by the “inaccurate” modification. Using these synthetic profiles, we then format the data for instruction fine-tuning, where each sample has four parts: System, Instruction, Input, and Output.
@@
// Inference
@@
       // The Diagnostic Module
@@ 5       $\{l\}\leftarrow f_{\text{diag}}(\mathcal{P}_{u}^{(ini)},\mathcal{C})$;
       // The Treatment Module
6       $\{r\}\leftarrow\text{Infer}(\mathcal{P}_{u}^{(ini)},\mathcal{C},\{l\})$
@@ 7       $\mathcal{O}\leftarrow\text{GenOpt}(\mathcal{P}_{u}^{(ini)},\mathcal{C},\{l\},\{r\})$;
8       $\tilde{\mathcal{P}}_{u}^{(ini)}\leftarrow\text{Refine}(\mathcal{P}_{u}^{(ini)},\mathcal{O})$ $\mathcal{C}$ $\leftarrow$ [ ];
@@ 9       $\mathcal{P}_{u}^{(ini)}$ $\leftarrow$ $\tilde{\mathcal{P}}_{u}^{(ini)}$;
This module is responsible for the complex reasoning and modification tasks. While the specialized diagnostic module excels at identifying what defects exist within a profile, understanding why these defects occur and determining how to address them effectively requires sophisticated common-sense reasoning and high-fidelity generative capabilities. Consequently, we employ a powerful general-purpose LLM, leveraging its superior analytical reasoning and text generation capabilities to carry out the following treatment process.
@@
Existing LLM-based user simulators are primarily integrated with static recommenders in page-by-page scenarios (zhang2024generative), which essentially represent a one-time recommendation result split into multiple pages. This fails to capture realistic scenarios where users continuously interact with the RSs, leading to the dynamic evolution of both user profiles and recommendation strategies through user feedback and interactions. To enable realistic multi-round interactions, we propose to incorporate our optimized user simulator DGDPO with sequential recommenders (SRs). Unlike traditional static models, SRs excel at capturing temporal dynamics, which makes them ideal for modeling the bidirectional evolution of user preferences and recommendation strategies over time.
@@
We conduct extensive experiments to answer five research questions [^1] Our code and datasets are available at https://github.com/hyllll/DGDPO. (RQ1) Can DGDPO improve the fidelity of user simulators? (RQ2) How do different profile update strategies and SRs impact the simulation fidelity in multi-round interaction scenarios? (RQ3) How do various components of DGDPO affect its performance? (RQ4) How do essential hyper-parameters affect DGDPO? (RQ5) How does DGDPO construct user profile with its optimization process?
@@
We use three datasets: ML-1M (harper2015movielens), Amazon Books, and Amazon Movies & TV (ni2019justifying). We filter users and items with less than 5 interactions (ML-1M) or less than 20 (Amazon datasets). Interactions are sorted chronologically, with max sequence lengths set to 200 for ML-1M and 50 for Amazon datasets. For each user, the last 10 items are for testing, the 11th most recent item is for validation, and the rest for training (i.e., $\mathcal{D}_{u}$). $\mathcal{D}_{u}$ is further split into $\mathcal{D}_{u}^{(ini)}$ and $\mathcal{D}_{u}^{(opt)}$ by an $\alpha:(1-\alpha)$ ratio.
@@
We compare DGDPO with five baselines for user profile construction in two categories. In traditional methods, KuaiSim (zhao2023kuaisim) is a transformer-based model for user profile construction. In LLM-based methods, Agent4Rec (zhang2024generative) utilizes a single-step prompt-based inference. LLM-US (zhang2025llm) generates explanations of user ratings as profile. RAH (shu2024rah) and AgentCF (zhang2024agentcf) employ self-reflection and collaborative reflection, respectively. For multi-round interactions, we integrate the user simulator with four SRs: FPMC (rendle2010factorizing), NARM (li2017neural), SASRec (kang2018self) and LRURec (yue2024linear).
@@
Following state-of-the-arts (shu2024rah; zhang2024generative), we randomly sample 1000 users from each dataset to serve as simulated users. Our diagnostic module is based on lightweight Qwen3-0.6B (yang2025qwen3), while the treatment module employs gpt-4o-mini via its API with a temperature of 0. All LLM-based baselines also utilize GPT-4o-mini as their backbone. For DGDPO, we set $N_{1}=4$, $\alpha=0.6$, and $\mathcal{X}=20$ for all datasets. We build upon Agent4Rec (zhang2024generative) as the foundation user simulator and conduct ten-round interactions for each user. For KuaiSim and SRs methods, we tune them by Optuna (optuna.org) with 50 trials and train for up to 100 epochs with an early stopping mechanism (sun2022daisyrec). All experiments are conducted on a V100 GPU.
@@
Following (luo2022mindsim; zhang2024generative), we adopt Precision (Prec), Recall, Accuracy (Acc), and F1 Score (F1) as the evaluation metrics to measure the consistency between the simulated and real users.
@@
Fig. 3 depicts the performance of DGDPO across ten interaction rounds on Books dataset under different profile update strategies [^2] Hereafter, we only show the results on Books due to space limitation, and similar trends can be noted on the other two datasets.. The results unveils that dynamic profile update in multi-round interactions can further improve the simualtion accuracy. (1) For ‘w/o Update’, the performance generally exhibits minimal fluctuations across multi-round interactions and is the worst among all strategies. With profile unchanged, ‘w/o Update’ fails to capture the evolution of user preference during multi-round interactions, thereby limiting the simulator’s long-term fidelity. (2)Update w/o GT’ exhibits performance divergence when interacting with different SRs. Specifically, with FPMC, the performance shows a declining trend. In contrast, with advanced deep learning-based SRs, we observe steady performance improvements. This occurs because advanced SRs rank ground truth items higher, and LLM-based user simulators, sensitive to item positions (hou2024large), are more likely to interact with them when they are prominently placed, resulting in more accurate simulation. (3)Update w/ GT’ performs the best among all methods. This is attributed to an idealized update strategy that always uses the ground truth item to update the user profile, regardless of the simulator’s actions. However, in practical applications, ground truth items are often unavailable, which limits the applicability of this strategy. Therefore, it underscores the importance of developing more advanced user simulators that closely replicate real user behaviors, enabling more effective and reliable updates to the user profile during long-term interactions.
@@
Fig. 5 illustrates the DGDPO optimization process with a case study from the Books dataset. A discrepancy (a highly-rated book being ignored by the simulator) triggers the diagnostic module to identify the profile as “Incomplete”. The treatment module then infers the missing preference for “Politics & Social Sciences” and updates the profile accordingly.